Hello,
a while back, I had submitted this problem:
simplify (5^2 3^2 7^1)^2
_______________
5^4
It was answered, but does not match any of the choices below:
a) 63 b) 3,969 c) 3,556,224 d) 4,064,256
_________ _________
625 625

Answer by Kodimalar(1)

The expression (5^2*3^2*7^1)^2/5^4 can be written as 5^4*3^4*7^2/5^4.
Cancelling 5^4 in the numerator and denominator, 3^4*7^2 = 81*49 = 3969
Answer: (b)